5. (10 points) Find the general solution of the following differential equations. 4y-12y+9y = 0 (0) = 2 y(0) = 5 6. (10 po
Answer #1


③ 44 - 12 y +9 y=0 y (6)=2, y(0) = 5 Let y=mx be a trial solution. Then the auxiliary equation is 4 m² - 12m +9 = 0 = (2m -y=emt y + y = et + cost-at Let be a trial solution of y + y =0. Then the auxiliary equation is m² + m²=0 m2 (m+1)=0 => m=0,0,